A sample of CO2 gas at 25 ºC has a volume of 15 L and exerts a pressuer of 15.6 psi. How many grams of gas are in the sample?




1
Expert's answer
2021-04-07T11:29:05-0400

We use Ideal Gas Law;

PV=nRT

Therefore; "n=\dfrac{PV}{RT}"

Since;

P=15.6 psi (changing it to atm we divide by 14.696) = "\dfrac{15.6psi}{14.696} = 1.0615atm"

V= 15L

T= (25 + 273) = 298K

R = 0.0821atm L mol-1 K-1

"n= \dfrac{1.0615atm x 15L}{0.0821Latm mol-1K-1 x 298K}" = 0.650806 moles


If moles ="\dfrac{mass}{MM}"

Mass = Moles x MM

Since MM of CO2 = 44.01g/mol


Mass = 0.0650806mol x 44.01g/mol

= 28.642g